OpenStudy (anonymous):

Eliminate the parameter. x = t - 3, y = t^2 + 5

OpenStudy (raffle_snaffle):

?

OpenStudy (anonymous):

@raffle_snaffle ??

OpenStudy (raffle_snaffle):

I don't understand what you are asking but I am sure there is someone else who knows better than me

OpenStudy (anonymous):

same^

OpenStudy (anonymous):

Solve for \(t\)

OpenStudy (anonymous):

It's clear that \(x\) has a one to one relationship with \(t\) and \(y\) doesn't, so you want to find \(t=g(x)\).

OpenStudy (anonymous):

x=3 and y= -sqrt5? @wio

OpenStudy (anonymous):

Then you find out:\[ y= [g(x)]^2+5 \]

OpenStudy (watchmath):

\(t=x+3\) \(y=(x+3)^2+5=x^2+6x+14\)

OpenStudy (anonymous):

@watchmath @wio Thanks guys, I think I get where you got that from!(: